# Language-guided Skill Learning with Temporal Variational Inference

### v1

@@ Abstract
We present an algorithm for skill discovery from expert demonstrations. The algorithm first utilizes Large Language Models (LLMs) to propose an initial segmentation of the trajectories. Following that, a hierarchical variational inference framework incorporates the LLM-generated segmentation information to discover reusable skills by merging trajectory segments. To further control the trade-off between compression and reusability, we introduce a novel auxiliary objective based on the Minimum Description Length principle that helps guide this skill discovery process. Our results demonstrate that agents equipped with our method are able to discover skills that help accelerate learning and outperform baseline skill learning approaches on new long-horizon tasks in BabyAI, a grid world navigation environment, as well as ALFRED, a household simulation environment.
@@ 1 Introduction
Figure 2: Overall framework of LAST. Step 1: given a dataset of expert demonstrations, we query an LLM (only using the goal and actions as input) for an initial segmentation and a language description for each segment. Step 2: Language-augmented temporal variational inference takes in multi-modal data as input to improve upon the segmentation by merging different subsequences into skills. Step 3: online hierarchical RL on new tasks leveraging the learned skills which can greatly shorten the task horizon and help the agent efficiently learn on new tasks.
@@
Specifically, we propose a novel algorithm that can discover reusable skills with common structure via temporal variational inference on expert trajectories, augmented with language annotations. Our contributions are: 1) We propose a method that leverages a pretrained LLM to generate an initial segmentation of the expert trajectories, which greatly decreases the search space for skill segmentation and provides additional learning signals for the following skill discovery steps. 2) We propose a hierarchical variational inference framework that can incorporate the generated language supervision and discover skills on top of initial trajectory segmentations. 3) We augment the temporal variational inference process with a novel auxiliary training objective following Minimum Description Length Principle, which further helps compression and discover reusable skills. 4) The proposed online hierarchical RL framework enables the agent to quickly adapt to new long-horizon tasks with the learned skills in BabyAI (Chevalier-Boisvert et al., 2019) and ALFRED (Shridhar et al., 2020a), the latter of which is a highly complex household simulation environment with multimodal inputs and long-horizon tasks.
@@ 2 Related Work
Recent studies have proposed using natural language to guide the discovery of reusable skills (Sharma et al., 2022) from unparsed demonstrations. However, they have assumed that the skills themselves must be represented with natural language . Although powerful in “realistic” domains, natural language as the choice of representation can become restrictive when the optimal skill set is hard to specify using natural language alone. In this paper, we use language (and language models) to guide the discovery of skills, while lifting the restriction. Other recent work simultaneously propose the domain model and learn libraries of action operators defined as code-based policies guided by LLMs (Wang et al., 2023; Liu et al., 2023; Wong et al., 2023) as a way to learn hierarchical action representations. However, these approaches require a planner and multiple rounds of interaction in the environment thereafter to iteratively verify the correctness and extensibility of the inferred skills.
@@ 4.1 Initial Segmentation with LLMs
The core of LAST is to segment trajectories of $T$ actions into $N$ subsequences. One could imagine directly learning the graphical model. However, there is an exponential number of ways to segment a sequence of $T$ actions into $N$ subsequences, which will make the learning difficult and prone to poor local optima. Besides, the posterior $p(\phi\mid\tau)$ will be difficult to approximate. To see this, note that, given a trajectory $\tau$, the variables $\beta_{t}$ are heavily correlated since switching skills at timestep $t$ affects the probability of switching skills at nearby timesteps. Hence, approximating the posterior with, e.g., a factorial distribution over the $\beta_{t}$ will be very inaccurate and so will be our skill extraction procedure.
To address this, we shall thus first perform a segmentation of each trajectory into subsequences by using an LLM. There is a tradeoff to achieve. On the one hand, having short subsequences will only mildly simplify the posterior distribution. On the other hand, having few, long subsequences will make the posterior easier to approximate but at the expense of a loss of capacity in our generative model, especially if the segmentation is inaccurate. As we wish to make few assumptions about the quality of the underlying LLM, we only use it to find short segments, i.e., 1 to 5 actions each.
@@
The output of this step is a set of enriched trajectories where, at each timestep $t$, we added variables $\bar{\beta}_{t}$ and $g_{t}$. $\bar{\beta}_{t}$ equals 0 when $a_{t-1}$ and $a_{t}$ belonged to the same segment, 1 otherwise. $g_{t}$ denotes the language description generated by the LLM (Navigate to kitchen step 1) for the segment that contains $a_{t}$ ($g_{t}$ remains the same for the whole segment). It provides the additional information that we can use to further improve the initial segmentation. We call the first variable $\bar{\beta}$, not $\beta$, because it is not the final segmentation of the trajectory.
@@ 4.2 Skill Discovery with Temporal Variational Inference
$J(\theta)\approx$ $\mathbb{E}_{q_{\theta}(\phi\mid\tau)}\sum_{t=1}^{T}\log{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\pi_{\theta}(a_{t}\mid o_{:t},k_{t},G)}$ (3)
@@
$\quad\>\>\>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}p_{\theta}(\beta_{t}\mid o_{:t},a_{:t-1},\beta_{:t-1},k_{:t-1},G)}]$
$+$ $\text{KL}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}q_{\theta}(k_{t}\mid o_{:T},a_{:T},\beta_{t},k_{t-1},g_{:T},G)}\|$
$\quad\>\>\>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}p_{\theta}(k_{t}\mid o_{:t},a_{:t-1},k_{:t-1},\beta_{:t},G)}]\Big{\}},$
@@ 4.3 Minimum Description Length for Skills
We focus on minimizing the first term $L(\mathcal{D}\mid\phi)$. Recall that we use $\phi=\{\beta_{t},k_{t}\}_{t=1}^{T}$ to encode the trajectory $\tau=\{G,o_{t},a_{t},\bar{\beta}_{t},g_{t}\}_{t=1}^{T}$. According to optimal code length theory (Cover & Thomas, 2001), the expected number of bits of code generated by $p(k)$ is: $\mathbb{E}_{k}[-\log p(k)]$. Applying the formulation defined in § 4.2, we can derive the following objective:
@@ 5.1 Experimental Setup
ALFRED (Shridhar et al., 2020a) is an complex environment based on the AI2-THOR (Kolve et al., 2017) simulator and is a domain where an agent is required to perform diverse household tasks following a goal described by natural language. The environment itself contains $100+$ floorplans and $100+$ objects to interact with. The observation space consists of the agent’s egocentric view (image) of the current environment. The action space consists of 12 discrete action types (e.g., move forward, turn right, toggle) and 82 discrete object types. At each time step the agent needs to choose an action type along with an object type to interact with the environment. For the ALFRED dataset, we follow the settings in Pashevich et al. (2021) where the training dataset consists of more than 20000 trajectories. Each trajectory includes a goal description and a sequence of observations and actions. Note that different from many previous work on ALFRED (especially the leaderboard methods) we do not use the step-by-step instructions provided by environment (neither during training nor testing), which gives us a standard goal-conditioned RL setting and is potentially more practical as instructions annotated by humans could be expensive. We want to highlight that the goal of this work is not to directly compare to systems on the ALFRED leaderboard in terms of overall performance, rather, we adopt ALFRED as a challenging testbed demonstrating our proposed skill discovery method.
@@
Table 1: Online (zero-shot) task success rate comparison on BabyAI (top) and ALFRED (bottom).
@@ Acknowledgement
This work was partially done while Haotian Fu, Pratyusha Sharma, and Elias Stengel-Eskin were interning at MSR. The authors would like to thank Sumana Basu, Ayush Jain, Yiding Jiang, Michael Littman, Vincent Micheli, Benjamin Spiegel, Tongzhou Wang, Tian Yun, Zilai Zeng, as well as other colleagues and friends at Microsoft Research and Brown University for discussions and helpful feedback.
Broader Impact
@@ Appendix A Example Input&Output of Initial Segmentation with LLM
Input: Goal:place a microwaved apple slice on top of the black table Action sequence: LookDown15, MoveAhead300, RotateRight90, MoveAhead100, LookDown15, PickupObject ButterKnife, SliceObject Apple, LookUp15, RotateLeft180, MoveAhead25, RotateRight90, LookDown15, PutObject DiningTable, LookUp15, RotateRight90, MoveAhead25, LookDown15, PickupObject Apple, LookUp15, RotateRight90, MoveAhead25, RotateRight90, MoveAhead150, RotateLeft90, MoveAhead250, RotateRight90, MoveAhead75, RotateLeft90, MoveAhead25, OpenObject Microwave, PutObject Microwave, CloseObject Microwave, ToggleObjectOn Microwave, ToggleObjectOff Microwave, OpenObject Microwave, PickupObject Apple, CloseObject Microwave, RotateLeft90, MoveAhead100, RotateLeft90, MoveAhead300, RotateRight90, MoveAhead75, PutObject SideTable. Output: Action alignment: ,
@@ Appendix D Temporal Variational Inference
$+\log\pi(a_{t}\mid o_{:t},k_{t},G)+\log p({\beta}_{t}\mid o_{:t},a_{:t-1},\beta_{:t-1},k_{:t-1},G)+\log p(k_{t}\mid o_{:t},a_{:t-1},k_{:t-1},\beta_{:t},G)\}-\log q(\beta_{:T},k_{:T}\mid\cdot)]$
@@
$=\mathbb{E}_{q_{\theta}(\phi\mid\tau)}\sum_{t=1}^{T}\log{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\pi_{\theta}(a_{t}\mid o_{:t},k_{t},G)}-\sum_{t=1}^{T}\Big{\{}\text{KL}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}q_{\theta}(\beta_{t}\mid o_{:T},a_{:T},\bar{\beta}_{t},g_{:T},G)}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}p_{\theta}(\beta_{t}\mid o_{:t},a_{:t-1},\beta_{:t-1},k_{:t-1},G)}]$
[…]

### v2

@@ Abstract
We present an algorithm for skill discovery from expert demonstrations. The algorithm first utilizes Large Language Models (LLMs) to propose an initial segmentation of the trajectories. Following that, a hierarchical variational inference framework incorporates the LLM-generated segmentation information to discover reusable skills by merging trajectory segments. To further control the trade-off between compression and reusability, we introduce a novel auxiliary objective based on the Minimum Description Length principle that helps guide this skill discovery process. Our results demonstrate that agents equipped with our method are able to discover skills that help accelerate learning and outperform baseline skill learning approaches on new long-horizon tasks in BabyAI, a grid world navigation environment, as well as ALFRED, a household simulation environment. [^1] Code and videos can be found at https://language-skill-discovery.github.io/.
@@ 1 Introduction
Figure 2: Overall framework of LAST. Step 1: given a dataset of expert demonstrations, we query an LLM (only using the goal and actions as input) for an initial segmentation and a language description for each segment. Step 2: temporal variational inference takes in multi-modal data as input to improve upon the segmentation by merging different subsequences into skills. Step 3: online hierarchical RL on new tasks leveraging the learned skills which can greatly shorten the task horizon and help the agent efficiently learn on new tasks.
@@
Specifically, we propose a novel algorithm that can effectively merge semantic priors from language models with temporal variational inference to discover skills. Our contributions are: 1) We propose a method that leverages a pretrained LLM to generate an initial segmentation of the expert trajectories, which greatly decreases the search space for skill segmentation and provides additional learning signals for the following skill discovery steps. 2) We propose a hierarchical variational inference framework that can incorporate the generated language supervision and discover skills on top of initial trajectory segmentations. 3) We augment the temporal variational inference process with a novel auxiliary training objective following Minimum Description Length Principle, which further helps compression and discover reusable skills. 4) The proposed online hierarchical RL framework enables the agent to quickly adapt to new long-horizon tasks with the learned skills in BabyAI (Chevalier-Boisvert et al., 2019) and ALFRED (Shridhar et al., 2020a), the latter of which is a highly complex household simulation environment with multimodal inputs and long-horizon tasks.
@@ 2 Related Work
The proposed approach improves over previous work in two ways: First, methods that use language/language model priors, inadvertently restrict the representation of the skills to natural language alone—this restriction doesn’t apply to our framework. Second, most methods that use variational inference to discover skills from demonstration do not utilize any external semantic guidance from language models, which makes the skill discovery procedure prone to overfitting to short-horizon plans and the model fails to learn skills that generalize well.
@@
Recent studies have proposed using natural language to guide the discovery of reusable skills (Sharma et al., 2022) from unparsed demonstrations. However, they have assumed that the skills themselves must be represented with natural language . Although powerful in “realistic” domains, natural language as the choice of representation can become restrictive when the optimal skill set is hard to specify using natural language alone. In this paper, we use language (and language models) to guide the discovery of skills, while lifting the restriction. BOSS (Zhang et al., 2023b) starts from a pretrained initial skill library and focuses on learning new skills through online interactions by chaining the existing skills which are guided by LLM directly. SPRINT (Zhang et al., 2023a) proposes to use LLMs to relabel robot trajectories and leverage offline reinforcement learning methods to do skill chaining. Other recent work simultaneously propose the domain model and learn libraries of action operators defined as code-based policies guided by LLMs (Wang et al., 2023; Liu et al., 2023; Wong et al., 2023) as a way to learn hierarchical action representations. However, these approaches require a planner and multiple rounds of interaction in the environment thereafter to iteratively verify the correctness and extensibility of the inferred skills.
@@ 4.1 Initial Segmentation with LLMs
The core of LAST is to segment trajectories of $T$ actions into $N$ subsequences. One could imagine directly learning the graphical model. However, there are an exponential number of ways to segment a sequence of $T$ actions into $N$ subsequences, which will make the learning difficult and prone to poor local optima. Besides, the posterior $p(\phi\mid\tau)$ will be difficult to approximate. To see this, note that, given a trajectory $\tau$, the variables $\beta_{t}$ are heavily correlated since switching skills at timestep $t$ affects the probability of switching skills at nearby timesteps. Hence, approximating the posterior with, e.g., a factorial distribution over the $\beta_{t}$ will be very inaccurate and so will be our skill extraction procedure.
To address this, we shall thus first perform a segmentation of each trajectory into subsequences by using an LLM. There is a tradeoff to achieve. On the one hand, having short subsequences will only mildly simplify the posterior distribution. On the other hand, having few, long subsequences will make the posterior easier to approximate but at the expense of a loss of capacity in our generative model, especially if the segmentation is inaccurate. As we wish to make few assumptions about the quality of the underlying LLM, we only use it to find short segments, i.e., 1 to 5 actions each. This constraint is introduced by modifying the prompt to the LLM, as shown in Appendix B: in the prompt, we tell the LLM that the number of actions assigned to each skill should not exceed 5 but should be larger than 1.
@@
The output of this step is a set of enriched trajectories where, at each timestep $t$, we added variables $\bar{\beta}_{t}$ and $g_{t}$. $\bar{\beta}_{t}$ equals 0 when $a_{t-1}$ and $a_{t}$ belonged to the same segment, 1 otherwise. $g_{t}$ denotes the language description generated by the LLM (Navigate to kitchen step 1) for the segment that contains $a_{t}$ ($g_{t}$ remains the same for the whole segment). The language annotations $g_{t}$ generated by the LLM are also an important conditioning information for our model to do the inference during the temporal variational inference phase. We call the first variable $\bar{\beta}$, not $\beta$, because it is not the final segmentation of the trajectory.
@@ 4.2 Skill Discovery with Temporal Variational Inference
$J(\theta)\approx$ $\mathbb{E}_{q_{\theta}(\phi\mid\tau)}\sum_{t=1}^{T}\log{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pi_{\theta}(a_{t}\mid o_{:t},k_{t},G)}$ (3)
@@
$\quad\>\>\>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}p_{\theta}(\beta_{t}\mid o_{:t},a_{:t-1},\beta_{:t-1},k_{:t-1},G)}]$
$+$ $\text{KL}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}q_{\theta}(k_{t}\mid o_{:T},a_{:T},\beta_{t},k_{t-1},g_{:T},G)}\|$
$\quad\>\>\>{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}p_{\theta}(k_{t}\mid o_{:t},a_{:t-1},k_{:t-1},\beta_{:t},G)}]\Big{\}},$
@@ 4.3 Minimum Description Length for Skills
We focus on minimizing the first term $L(\mathcal{D}\mid\phi)$. Recall that we use $\phi=\{\beta_{t},k_{t}\}_{t=1}^{T}$ to encode the trajectory $\tau=\{G,o_{t},a_{t},\bar{\beta}_{t},g_{t}\}_{t=1}^{T}$. According to optimal code length theory (Cover & Thomas, 2001), the expected number of bits of code generated by $p(k)$ is: $\mathbb{E}_{k}[-\log p(k)]$. Applying the formulation defined in § 4.2, we can derive the following objective (see Appendix F for detailed derivations):
@@ 5.1 Experimental Setup
ALFRED (Shridhar et al., 2020a) is a complex environment based on the AI2-THOR (Kolve et al., 2017) simulator and is a domain where an agent is required to perform diverse household tasks following a goal described by natural language. The environment itself contains $100+$ floorplans and $100+$ objects to interact with. The observation space consists of the agent’s egocentric view (image) of the current environment. The action space consists of 12 discrete action types (e.g., move forward, turn right, toggle) and 82 discrete object types. At each time step the agent needs to choose an action type along with an object type to interact with the environment. For the ALFRED dataset, we follow the settings in Pashevich et al. (2021) where the training dataset consists of more than 20000 trajectories. Each trajectory includes a goal description and a sequence of observations and actions. Note that different from many previous work on ALFRED (especially the leaderboard methods) we do not use the step-by-step instructions provided by environment (neither during training nor testing), which gives us a standard goal-conditioned RL setting and is potentially more practical as instructions annotated by humans could be expensive. We want to highlight that the goal of this work is not to directly compare to systems on the ALFRED leaderboard in terms of overall performance, rather, we adopt ALFRED as a challenging testbed demonstrating our proposed skill discovery method.
@@
Table 1: Online (zero-shot) task success rate comparison on BabyAI (top) and ALFRED (bottom). Note that the subtasks in ALFRED have short horizons (i.e., subgoals of each individual task which requires significantly fewer steps to reach). In comparison, in the downstream task learning setting, the agent is tested on six individual tasks with the full length.
@@ 6 Discussion
The proposed initial segmentation approach requires agents equipped with discrete action tokens. This is one limitation of the proposed method. However, we would like to point out that many practical applications have the same settings (e.g., policies that use API calls or tools, AI assistants that help with typical performance in a program or OS, gaming settings such as Atari where actions tend to be discretized). Besides, we should also note that the core idea of initial segmentation is to leverage the commonsense knowledge from foundation models to perform the initial segmentation. At the moment, we are limited to using Large Language Models (hence using the text description of the actions). But it is also possible to use large video foundation models to improve the initial segmentation and extend our work to other domains in the future.
@@ Acknowledgement
This work was partially done while Haotian Fu, Pratyusha Sharma, and Elias Stengel-Eskin were interning at MSR. The authors would like to thank Sumana Basu, Ayush Jain, Yiding Jiang, Michael Littman, Vincent Micheli, Benjamin Spiegel, Tongzhou Wang, Tian Yun, Zilai Zeng, ICML anonymous reviewers, as well as other colleagues and friends at Microsoft Research and Brown University for discussions and helpful feedback. This work was supported in part by NSF grant #1955361, CAREER award #1844960 to Konidaris, and ONR grant #N00014-22-1-2592. And Nicolas Le Roux was supported by a CIFAR AI chair.
Impact Statement
@@ Appendix A Example Input&Output of Initial Segmentation with LLM
Input: Goal:place a microwaved apple slice on top of the black table Action sequence: LookDown15, MoveAhead300, RotateRight90, MoveAhead100, LookDown15, PickupObject ButterKnife, SliceObject Apple, LookUp15, RotateLeft180, MoveAhead25, RotateRight90, LookDown15, PutObject DiningTable, LookUp15, RotateRight90, MoveAhead25, LookDown15, PickupObject Apple, LookUp15, RotateRight90, MoveAhead25, RotateRight90, MoveAhead150, RotateLeft90, MoveAhead250, RotateRight90, MoveAhead75, RotateLeft90, MoveAhead25, OpenObject Microwave, PutObject Microwave, CloseObject Microwave, ToggleObjectOn Microwave, ToggleObjectOff Microwave, OpenObject Microwave, PickupObject Apple, CloseObject Microwave, RotateLeft90, MoveAhead100, RotateLeft90, MoveAhead300, RotateRight90, MoveAhead75, PutObject SideTable. Output: Action alignment: [ { index: 0, summary action str: navigate to kitchen step 1, robot actions str: LookDown 15, MoveAhead300 }, { index: 1, summary action str: navigate to kitchen step 2, robot actions str: RotateRight90, MoveAhead100, LookDown15 }, { index: 2, summary action str: ”pick up butterknife and slice apple”, robot actions str: ”PickupObject ButterKnife, SliceObject Apple” }, { index: 3, summary action str: ”return butterknife to table”, robot actions str: ”LookUp15, RotateLeft180, MoveAhead25, RotateRight90, LookDown15, PutObject DiningTable” }, { index: 4, summary action str: ”pick up apple slice”, robot actions str: ”LookUp15, RotateRight90, MoveAhead25, LookDown15, PickupObject Apple” }, { index: 5, summary action str: ”navigate to microwave step 1”, robot actions str: ”LookUp15, RotateRight90, MoveAhead25, RotateRight90, MoveAhead150” }, { index: 6, summary action str: ”navigate to microwave step 2”, robot actions str: ”RotateLeft90, MoveAhead250, RotateRight90, MoveAhead75” }, { index: 7, summary action str: ”navigate to microwave step 3”, robot actions str: ”RotateLeft90, MoveAhead25” }, { index: 8, summary action str: ”place apple slice in microwave and turn on”, robot actions str: ”OpenObject Microwave, PutObject Microwave, CloseObject Microwave, ToggleObjectOn Microwave, ToggleObjectOff Microwave” }, { index: 9, summary action str: ”retrieve microwaved apple slice”, robot actions str: ”OpenObject Microwave, PickupObject Apple, CloseObject Microwave” }, { index: 10, summary action str: ”navigate to black table step 1”, robot actions str: ”RotateLeft90, MoveAhead100, RotateLeft90, MoveAhead300” }, { index: 11, summary action str: ”navigate to black table step 2”, robot actions str: ”RotateRight90, MoveAhead75” }, { index: 12, summary action str: ”place microwaved apple slice on black table”, robot actions str: ”PutObject SideTable” } ],
@@ Appendix D Temporal Variational Inference
$+\log\pi(a_{t}\mid o_{:t},k_{t},G)+\log p({\beta}_{t}\mid o_{:t},a_{:t-1},\beta_{:t-1},k_{:t-1},G)+\log p(k_{t}\mid o_{:t},a_{:t-1},k_{:t-1},\beta_{:t},G)\}$
$-\log q(\beta_{:T},k_{:T}\mid\cdot)]$
@@
$=\mathbb{E}_{q_{\theta}(\phi\mid\tau)}\sum_{t=1}^{T}\log{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pi_{\theta}(a_{t}\mid o_{:t},k_{t},G)}-\sum_{t=1}^{T}\Big{\{}\text{KL}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}q_{\theta}(\beta_{t}\mid o_{:T},a_{:T},\bar{\beta}_{t},g_{:T},G)}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}p_{\theta}(\beta_{t}\mid o_{:t},a_{:t-1},\beta_{:t-1},k_{:t-1},G)}]$
$+\text{KL}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}q_{\theta}(k_{t}\mid o_{:T},a_{:T},\beta_{t},k_{t-1},g_{:T},G)}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}p_{\theta}(k_{t}\mid o_{:t},a_{:t-1},k_{:t-1},\beta_{:t},G)}]\Big{\}}$